OpenStudy (anonymous):

All right, medal for best and most creative answer!!! does not have to make sense!!! What is 1+1?

OpenStudy (anonymous):

fish*batman/superman+wonderwoman^hulk

OpenStudy (anonymous):

or could it be deer^hog-bat+dracula*viper

OpenStudy (anonymous):

Hmm... best if it has some sort of correllation to 1+1... such as: 1+1=infinity because 1 universe plus 1 black hole = the infinite cosmos...one black hole leads to an alternate universe, which has many black holes of its own... which all lead to seperate alternate universes, or alternate universes and multiverses that cross over each other... which create an infinite number of universes. Each single universe is infinite, and infinity+infinity is infinite, so 1+1=infinity. Of course, without that first universe, there couldn't be a black hole, and therefore there can't be that web of universes creating the great, infinite cosmos of the multiverse.

OpenStudy (anonymous):

\[pepsi \superman[\coke{snake0125}657 & thesselonians \\]

OpenStudy (anonymous):

now that's starting to look more like a computer virus than the answer to 1+1...

OpenStudy (anonymous):

lol ok i got this

OpenStudy (anonymous):

nvm i aint got it im stealing yours

OpenStudy (anonymous):

1+1=infinity because 1 universe plus 1 black hole = the infinite cosmos...one black hole leads to an alternate universe, which has many black holes of its own... which all lead to seperate alternate universes, or alternate universes and multiverses that cross over each other... which create an infinite number of universes. Each single universe is infinite, and infinity+infinity is infinite, so 1+1=infinity. Of course, without that first universe, there couldn't be a black hole, and therefore there can't be that web of universes creating the great, infinite cosmos of the multiverse.

OpenStudy (anonymous):

hey, no way. i will report you for stealing my ingenuity.

OpenStudy (anonymous):

lol

OpenStudy (mary.rojas):

1+1= peace love happiness and cute little kitties that rule the world of course duhhhhhh! lol :)

OpenStudy (anonymous):

MY INGENUITY. MY BRAINPOWER IS TOO VALUABLE TO BE STOLEN. NO MEDAL FOR YOU. @papsmurf

OpenStudy (anonymous):

hi @mary.rojas glad to have you back in our creative circle.

OpenStudy (mary.rojas):

hey :)

OpenStudy (anonymous):

lol ok @vipergirl81 chill 1+1= faith hope and happiness but the best is love

OpenStudy (anonymous):

Hmm... I'd say I'm more into infinite cosmos stuff.

OpenStudy (mary.rojas):

you forgot cute kitties that rule the world :/ !!! lol

OpenStudy (anonymous):

But Cute kittens are TO BE WORSHIPPED.

OpenStudy (mary.rojas):

:)

OpenStudy (anonymous):

1+1=faith hope happiness and love that the little kittens have that help them rule the world and all of the cosmos and too infinite and beyond

OpenStudy (anonymous):

Yeah, but that still has nothing to do with 1+1...

OpenStudy (anonymous):

holy crap im missing the big picture

OpenStudy (wolf1728):

1 + 1 = 11 (silly grade school reasoning) 1 + 1 = 10 (binary arithmetic)

OpenStudy (mary.rojas):

no 1+1=1+1 lol @wolf1728

OpenStudy (anonymous):

YEAH, @wolf1728, that's good. but... more creative grey matter! our brain will become dusty without a stretch or two! Think: borderline insane yet still kindsorta makes sense.

OpenStudy (mary.rojas):

close this question before someone gets you in trouble.

OpenStudy (wolf1728):

hmmmmmm ... must start thinking more insanely

OpenStudy (anonymous):

1 + 1 is Beyonce's song...it really is :)